Academic Employment in Mathematics

AMS Standard Cover Sheet

This cover sheet is provided as an aid to departments in processing job applications. It should be included with your other application material. Please print or type. Do not send this form to the AMS.

Last (Family) Name:  \underline{Stein}
First Name or Initial:  \underline{William}
Middle Name or Initial: \underline{A}

Address through June 2000:

Home Phone

2041 Francisco Street, Apt. 14

\underline{(510) 883-9938}

Berkeley, CA 94709 e-mail address \underline{USA}

                                                                     was@math.berkeley.edu

Current Institutional Affiliation:

Work Phone

University of California at Berkeley, Department of Mathematics

\underline{(510) 643-0893}

Highest Degree and Source \underline{B.S.} Year of Ph.D. (optional) Ph.D. Advisor: \underline{Hendrik W. Lenstra} If the Ph.D. is not presently held, date on which you expect to receive it: \underline{May 2000}

Indicate the mathematical subject area(s) in which you have done research using, if applicable, the 1991 Mathematics Subject Classification. If listing more than one number, list first the one number which best describes your current primary interest.

Primary Interest \underline{11} Secondary Interests (optional) \underline{22}

Give a very brief synopsis of your current research interests in the box below (e.g. finite group actions on four-manifolds). Avoid special mathematical symbols.

Galois representations, modular forms, explicit approaches to modular abelian varieties.

Most recent position held, if any, post Ph.D. University or Company

Position Title  Dates

Indicate the position for which you are applying and position posting code, if applicable \underline{Postdoc} If unsuccessful for this position, would you like to be considered for a temporary position? X Yes No If yes, please check the appropriate boxes.

X Postdoctoral Position X 2+ Year Position X 1 Year Position

List the names and affiliations of up to four individuals who will provide letters of recommendation if asked. Mark the box provided for each individual whom you have already asked to send a letter.

1. \underline{Kenneth A. Ribet, UC Berkeley, ribet@math.berkeley.edu}

2. \underline{Hendrik W. Lenstra, UC Berkeley, hwl@math.berkeley.edu}

3. \underline{Barry C. Mazur, Harvard University, mazur@abel.math.harvard.edu}

4. \underline{Robert F. Coleman, UC Berkeley, coleman@math.berkeley.edu}